Review problem. One electron collides elastically with a second electron initially at rest. After the collision, the radii of their trajectories are 1.00 cm and 2.40 cm. The trajectories are perpendicular to a uniform magnetic field of magnitude 0.044 0 T. Determine the energy (in keV) of the incident electron.


1
Expert's answer
2021-02-03T16:15:46-0500

The velocity of the incident electron after the interaction can be found from the radius of its trajectory:


"v=\frac{eBR}{m}."

Its kinetic energy is


"E_K=\frac12 mv^2=\frac12 \frac{(eBR)^2}{m}."

The energy in keV will be


"E_{K\text{(kEV)}}=\frac{E_K}{1000e}=\frac{eB^2R^2}{2000m}=5.51\cdot10^{-22}\text{ keV}."